OpenStudy (anonymous):

need desperate help Identify the open intervals on which the function is increasing or decreasing. (Select all that apply.) h(x) = 3x − x3 Increasing: 1 (−, −1) (−1, 1) (1, ) none of these . Decreasing: 2 (−, −1) (−1, 1) (1, ) none of these .

OpenStudy (anonymous):

\[h(x)=3x-x^3\] take derivative of h \[h'(x)=3-3x^2\] Find criticle point by setting it equals to 0 \[0=3-3x^2\] \[3=3x^2\] \[1=x^2\] \[x=\pm 1\]

OpenStudy (anonymous):

We will have to check at a number left of -1, between -1 and 1 and number right of 1

OpenStudy (anonymous):

sorry i wrote the wrong info; please read the corrrected one. I posted it as a new question PLESE LOOK AT THE ONE THE NEW ONE I POSTEd

OpenStudy (anonymous):

sorry i wrote the wrong info; please read the corrrected one. I posted it as a new question PLESE LOOK AT THE ONE THE NEW ONE I POSTEd

OpenStudy (anonymous):

sorry i wrote the wrong info; please read the corrrected one. I posted it as a new question PLESE LOOK AT THE ONE THE NEW ONE I POSTEd

OpenStudy (anonymous):

Differentiate the function. Solve for the zeros of the derivative. Draw a number line with a verticle line at each of the zeros. These lines determine the intervals where the derivative of the fuction changes sign. Pick a number (easy one) that is in the interval of interest. Evaluate the derivative at this point. Is it positive or negative? If it is positive then the function is increasing in this interval. If it is decreasing then the function is decreasing in this interval. I usually pick the interval furthest on the left and evaluate this one. The signs will alternatley change as you enter each new interval from the left.